OpenStudy (anonymous):

find the logarithm of each number. log5..sqrt5=?

OpenStudy (anonymous):

ok, write sqrt 5 as a power of 5. \[\log_{5}5^{1/2}=\frac{1}{2}\log_{5}5=\frac{1}{2}*1=\frac{1}{2} \] the first two steps are common, import laws of logs that you should commit to memory, probably

OpenStudy (anonymous):

thank you ur a genius

OpenStudy (anonymous):

woops typo that should say "important". and thank you!

OpenStudy (anonymous):

using scientific notation find the logarithm of each number i dnt know how to do this at all two example log 12 and log 0.038